Another representation, in polar coordinates, of the point (1,5pi/6) is (___, 11pi/6)


1
Expert's answer
2021-12-03T06:04:43-0500

When graphing a polar coordinate with a negative radius, we move from the pole in the direction opposite the given positive angle (on the same line as the given angle but in the direction opposite to the angle from the pole). In our case, another representation of the point "(1,\frac{5\pi}6)" is "(-1,\pi+\frac{5\pi}6)=(-1, \frac{11\pi}6)".